A tank initially contains 60 gal of brine in which there is dissolved 8 lb of salt. Brine containing 2 lb of dissolved salt per gallon flows into the tank at the rate of 5 gal/min. The mixture is kept uniform by stirring, and the stirred mixture simultaneously flows out at the slower rate of 3 gal/min. How much salt is in the tank at any time?


1
Expert's answer
2021-01-19T02:31:59-0500

Let x(t) represent the amount of salt in the tank at time t, where t is given in minutes.

The flow rate in is = 5 gal/min and the flow rate out is = 3 gal/min. So the Volume in the tank at the time t is

V(t) = 60 + (5 - 3)t = 60 + 2t

in gallons.

The concentration of salt in the tank (and hence of the solution flowing out of the tank) is then

c0(t) = x(t) / V(t) = x(t) / (60 + 2t)

Since the incoming concentration is ci = 2 lb/gal, the net rate of change of the salt is

dx/dt = rici - r0c0

dx/dt = 2 • 5 - (3 • x(t)) / (60 + 2t)

10 = x'(t) + (3 • x(t)) / (60 + 2t)

which is linear differential equation, where p(t) = 3 / (60 + 2t) and q(t) = 10, so that an integration factor is

m(t) = e^(-"\smallint"p(t)dt) = e^(-3"\smallint" dt/(60 + 2t)) = (60 + 2t)-3

Hence we have

x(t) / (60 + 2t)3 = "\smallint" 10dt / (60 + 2t)3

x(t) / (60 + 2t)3 = 5 / (60 + 2t)2 - C

x(t) = 60 + 2t - C(60 + 2t)3

Since x(0) = 8 then

8 = 60 + 0 - C(60 + 0)3

C = 13/54 000

Hence

x(t) = 60 + 2t - 13(60 + 2t)3/54 000